OpenStudy (anonymous):

If 2.891g of Mg Cl2 is dissolved in enough water to make 300.0mL of solution what is the molarity of the magnesium chloride solution? The molecular mass of magnesium chloride 95.3g/mol and the density of the solution is 1.22g/mol.

OpenStudy (aaronq):

Find the moles of magnesium chloride, divide by volume of solution. \(\sf moles=\dfrac{mass}{Molar~Mass}\) \(\sf Molarity=\dfrac{moles~of~solute}{L~of~solution}\)

OpenStudy (anonymous):

Its asking for the molarity of magnesium chloride within the solution.

OpenStudy (anonymous):

these are the possible answers to the problem .5505 Mol 5.782 Mol 5.782 x 10^-3 Mol 6.703 x 10^-2 Mol Also just noticed but is supposed to be 500mL not 300 mL.

OpenStudy (aaronq):

what i posted above is how you find the molarity of MgCl2 in the solution.

OpenStudy (aaronq):

i got 0.06067 M

OpenStudy (aaronq):

it's pretty much D, 0.06067 M = 6.07*10^-2 M that digit is likely an error

OpenStudy (anonymous):

what are you dividing? I did 500/1.22 and got 409.83 moles of solution. I did 2.891/95.3 and got 0.303357 moles of MgCl2 I did 500/95.3 and got 5.2465

OpenStudy (aaronq):

"density of the solution is 1.22g/mol." this is wrong mol is not a unit of volume, its a typographical error and they mean mL, 1.22 g/mL you dont need the density to solve the question though. i found moles, first \(\sf moles=\dfrac{2.891~ g}{95.3~ g/mol}\) then divided by the volume of solution (in liters) \(\sf Molarity=\dfrac{(\dfrac{2.891~ g}{95.3~ g/mol})} {0.5~L}\)

OpenStudy (anonymous):

ok I got it now. thank you for the explaination.